OpenStudy (anonymous):

Translate into a mathematical statement. " A number increased by 7 is 36"?

OpenStudy (anonymous):

n+7=36 or 7+n=36

OpenStudy (anonymous):

THANKS.

OpenStudy (anonymous):

How about, "Seven times a number is subtracted from15. The result is 42

OpenStudy (mertsj):

15-7n=42

OpenStudy (anonymous):

15-7n=42

OpenStudy (anonymous):

Nice, thanks again

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

anytime;)